OpenStudy (anonymous):

Which decimals in the list are repeating decimals? 1/3,3/5,4/15,5/6,4/1000,13/4

OpenStudy (anonymous):

A repeating decimal looks like this:\[\bf \frac{ 2 }{ 3 }=0.666666666...\]A non-repeating/terminating decimal looks like this:\[\bf \frac{ 1 }{ 10 }=0.1\]

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

So have you tried it yourself?

OpenStudy (anonymous):

yea 4 times

OpenStudy (anonymous):

K let's do a few:\[\bf \frac{ 1 }{ 3 }=0.33333...\]\[\bf \frac{3}{5}=0.6\]Which one is the repeating one from these two?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

So that's the first one that is repeating. Now let's look at the next two:\[\bf \frac{4}{15}=0.26666...\]\[\bf \frac{5}{6}=0.833333...\]Which is repeating from here? @ryanw775